A girl wants to buy her mother a birthday gift which is worth R900. She has only R400 and the birthday is in 10 days time. She decided to sell sweets to raise more money. On the first day she made a profit of 50 cents, on second day her profit was R1, on the third day her profit was R2, and so on until the 10th day.
1.2.1 Determine the total profit that the girl made.
1.2.2 Is she going to be able to afford the gift? Explain.

Answer by stanbon(75887)

A girl wants to buy her mother a birthday gift which is worth R900. She has only R400 and the birthday is in 10 days time. She decided to sell sweets to raise more money. On the first day she made a profit of 50 cents, on second day her profit was R1, on the third day her profit was R2, and so on until the 10th day.
1.2.1 Determine the total profit that the girl made.
sequence: (1/2),1,2,4,8,16,32,64,128,256
Sum = a[1-r^10)/(1-r)] = (1/2) = (1/2)[(1-2^10]= (1/2)[1023] = $511.50
--------
1.2.2 Is she going to be able to afford the gift? Explain.
Yes. She has 400 and is going to receive 511.50
================
Cheers,
Stan H.
==================
Cheers,
Stan H.